OpenStudy (anonymous):

calculus help: integrate (x+4)/x^2+2x+5

OpenStudy (anonymous):

I got stuck at (x+4)/((x+1)^2+4)

zepdrix (zepdrix):

See how you have an (x+4) in the numerator? I think you want to try and get an (x+4)^2 in the denominator. It will make a substitution easier to accomplish. Looks like we'll apply a trig sub after you get that sorted out.

OpenStudy (psymon):

I was also thinking of doing -4 + 4 in the numerator and breaking it into two integrals.

OpenStudy (psymon):

oops.

OpenStudy (psymon):

nvm x_x

zepdrix (zepdrix):

yah looks like that first one was gonna give us a lil trouble :c

OpenStudy (psymon):

I just noticed, haha. Yeah, trig-sub I like then.

zepdrix (zepdrix):

Oh we can't do an (x+4)^2 in the denominator can we? :P derp! my bad.

OpenStudy (psymon):

We can do trig-sub even though we don't have a square root, right?

zepdrix (zepdrix):

\[\large \int\limits \frac{x+4}{(x+1)^2+4}dx\] Yah we don't need a square root :) Hopefully this will work out.. hmm Making the substitution:\[\large x+1=2\tan \theta\]

zepdrix (zepdrix):

\[\large \int\limits \frac{(x+1)+3}{(x+1)^2+4}dx\qquad=\qquad \int\limits\frac{(2\tan \theta)+3}{4\tan^2\theta+4}dx\]Factoring out a 4 in the denominator,\[\large \frac{1}{4}\int\limits \frac{2\tan \theta+3}{\tan^2\theta+1}dx \qquad=\qquad \frac{1}{4}\int\limits\frac{2\tan \theta+3}{\sec^2\theta}dx\]Then we still need to deal with the dx. \[\large x+1=2\tan \theta \qquad\to\qquad dx=2\sec^2\theta\;d \theta\] Giving us,\[\large \frac{1}{4}\int\limits\limits\frac{2\tan \theta+3}{\cancel{\sec^2\theta}}(2\cancel{\sec^2\theta}\;d \theta)\]

zepdrix (zepdrix):

Ok I'll simmer down, I don't wanna do all it and take the fun away from ya :O Do you see where I'm going with this though? Stuck on any step in there?

OpenStudy (psymon):

Hmm...Lol, didn't even need to do what I'd have done. Yay for more than one way to do the trig sub.

OpenStudy (anonymous):

\[\int\frac{x+4}{x^2+2x+5}~dx=\frac{1}{2}\left(\int\frac{2x+2}{x^2+2x+5}~dx\right)+\int\frac{3}{(x+1)^2+4}~dx\] Not much different from what the others suggested, since it involves an algebraic and trig substitution, but the trig sub would be easier on the eyes.